Suppose that permanent income of consumption is calculated as the average of income over the past 5 years,that is YP=1/5(Y+Y-1+Y-2+Y-3+Y-4.Suppose that consumption is given by C=0.9YP.
a.If you have earned GH 20000 per year for the past 10 years.What is your permanent income
1
Expert's answer
2021-09-26T20:14:08-0400

"YP \ = \frac{1}{5}(Y + Y_1 + Y_2 + Y_3+ Y_4)"


I)

"YP = \frac{1}{5} \times (20000 \times 5)"

"YP = 20000"